# CityEQA: A Hierarchical LLM Agent on Embodied Question Answering Benchmark in City Space

### v2

@@ 4.3 Study on Collector Module
It is clear that the Collector significantly impacts outcomes. As Collector steps increase, NA decreases and QAA increases, suggesting that the Collector aids the agent in getting closer to targets and achieving accurate answers. However, there is a noticeable limit to QAA improvement; at Step 10, QAA is slightly lower than at Step 9. This may be due to the Collector’s poor judgment regarding action magnitude, resulting in "over-adjustment" of the observation and degrading visual input quality.
@@ 5 Related Works
5.1 EQA Datasets
Early research on using language to guide perception from visual input is known as Visual Question Answering (VQA) Ishmam et al. (2024); Guo et al. (2023). VQA tasks require agents to answer questions based solely on provided visual information (images or videos) Chandrasegaran et al. (2024). In contrast, Embodied Question Answering (EQA) involves agents actively navigating within an environment to seek visual inputs and enhance answer reliability Das et al. (2018). Due to cost and hardware limitations, several virtual indoor simulators have been developed for EQA tasks Liu et al. (2024a), resulting in indoor-focused datasets such as EQA-v1 Das et al. (2018) and MT-EQA Yu et al. (2019). Recently, urban environment simulators like EmbodiedCity Gao et al. (2024), CityNav Lee et al. (2024), and AerialVLN Liu et al. (2023) have emerged, though they mainly focus on navigation. EmbodiedCity provides an urban EQA dataset, but it functions more like VQA, as shown in Table 1. Moreover, due to the limited generalization capabilities of models at the time, only simple questions about basic attributes of objects were considered in these indoor datasetsRen et al. (2024). However, with the continuous improvement in the understanding and reasoning capabilities of pre-trained MM-LLMs for visual inputs, several open-ended EQA datasets have recently been released, such as K-EQA Tan et al. (2023) and OpenEQA Majumdar et al. (2024).
In comparison, this paper is the first to study the EQA tasks in city space and introduces the benchmark CityEQA-EC —- a high-quality dataset featuring diverse, open-vocabulary questions.
@@ 5.2 LLMs-driven Embodied Agents
The indoor EQA tasks mainly involve exploration and answer generation sub-tasks Ren et al. (2024). In early workDuan et al. (2022); Das et al. (2018); Lu et al. (2019), the two sub-tasks are mainly addressed by building and fine-tuning various deep neural networks. Recently, researchers attempt to utilize pre-trained LLMs to solve EQA tasks without any additional fine-tuningMu et al. (2024); Xiang et al. (2024); Huang et al. (2024). OpenEQA employed a Frontier-Based Exploration (FBE) strategy for indoor environment exploration and tested the performance of various MM-LLMs on the answer generation Majumdar et al. (2024). Besides, MM-LLMs was also used to determine which room to explore in indoor environment based their commonsense reasoning capabilities Yin et al. (2025).
These agents, however, cannot be directly used for CityEQA tasks. Unlike indoor spaces, which are confined and divided into rooms, city spaces are vast and open. Agents in cities must navigate using landmarks and spatial relationships for long-term exploration Zeng et al. (2024); Liu et al. (2024b). The proposed PMA addresses this by breaking down and planning for long-horizon CityEQA tasks, using large models across multiple modules to effectively handle open-ended questions and unseen environments.
@@ A.1 Dataset Collection and Validation
Any tasks identified as problematic were removed.
@@ A.3 Experiments Details
  • Blind Agents. We choose GPT-4 Achiam et al. (2023) and Qwen2.5 Yang et al. (2024) to answer questions as blind agents.
  • LLM-VQA. We choose GPT-4o and Qwen-2.5 to perform VQA tasks as LLM-VQA agents.
  • FBE Agent. Instead of utilizing landmarks and spatial relationships, it identifies the frontiers between explored and unexplored regions, samples one as the navigation point, and employs the A* algorithm to find a path. We also limit the path length to 10 meters at each step, consistent with the setting of the Navigator in the PMA.
@@
A Case of Human Agent
In Figure 10, we provide a case to illustrate why the performance of H_EQA is superior to that of H_VQA. The given question is "What is the color of the car next to the red car?" The ground truth answer is "Black". H_VQA was provided with the RGB image on the left for question answering. However, in this image, due to the influence of outdoor lighting, the originally black car appears gray, thus H_VQA provided an incorrect answer. In contrast, H_EQA can actively adjust the observation pose, observing the side of the car to reduce the impact of the lighting, and thereby providing the correct answer.
Figure 10: The images obtained by H-VQA and H-EQA.
Actions of Collector
The statistics of various actions taken by Collector are shown in Figure 11.
@@ A Case of Human Agent
Cases of Collector
We present two cases to illustrate the effect of the collector. In the first case, as shown in Figure 12, since the shop with black signboard was discovered too early in the Exploration stage, the starting pose of the collector was far from the target pose. Even after moving 10 steps promptly, it still failed to recognize the text on the black signboard. In the second case, as shown in Figure 13, in Step 1, the yellow signboard that the collector needed to recognize was on the left side of the picture and seemed not to be fully displayed. At this time, the collector took the TurnLeft action, thus observing the entire yellow signboard in Step 2 and easily providing the correct answer.
Figure 12: The failure case for collection.
Figure 13: The successful case for collection.

### v3

@@ 4.4 Effectiveness of Collector Module
The Collector significantly affects outcomes: as steps increase, NE decreases and QAA rises, helping the agent approach targets and improve accuracy. However, QAA plateaus, with Step 10 slightly lower than Step 9, possibly due to "over-adjustment" degrading visual input quality.
@@ 5 Related Works
5.1 QA and EQA
Early research on using language to guide perception from given input is known as Question Answering (QA), such as Visual QA (VQA) Ishmam et al. (2024) and 3DQA Zhan et al. (2025). These QA tasks require agents to answer questions based solely on provided information (images or cloud points) Chandrasegaran et al. (2024). In contrast, EQA involves agents actively exploring within an environment to seek visual inputs and enhance answer reliability Das et al. (2018). Due to cost and hardware limitations, several virtual indoor simulators have been developed for EQA tasks Liu et al. (2024a), resulting in indoor-focused datasets such as EQA-v1 Das et al. (2018) and MT-EQA Yu et al. (2019). However, although there are already several QA task datasets for outdoor environments, such as City-3DQA Sun et al. (2024) and Open3DVQA Zhan et al. (2025), EQA tasks have yet to be extended to outdoor settings, as shown in Table 1.
Recently, urban environment simulators like EmbodiedCity Gao et al. (2024), CityNav Lee et al. (2024), and AerialVLN Liu et al. (2023) have emerged, though they mainly focus on navigation task. EmbodiedCity provides an urban EQA dataset, but it functions more like VQA and ignores the active perception. Moreover, due to the limited generalization capabilities of models at the time, only simple questions about basic attributes of objects were considered in these indoor datasets Ren et al. (2024). However, with the continuous improvement in the understanding and reasoning capabilities of pre-trained VLMs for visual inputs, several open-ended EQA datasets have recently been released, such as Express-bench Jiang et al. (2025) and OpenEQA Majumdar et al. (2024). In comparison, this paper is the first to study the EQA tasks in city space and introduces the benchmark CityEQA-EC — a high-quality dataset featuring diverse, open-vocabulary questions.
@@ 5.2 LLMs-driven Embodied Agents
The indoor EQA tasks mainly involve exploration and answer generation sub-tasks Ren et al. (2024). In early work Duan et al. (2022); Das et al. (2018); Lu et al. (2019), the two sub-tasks are mainly addressed by building and fine-tuning various deep neural networks. Recently, researchers attempt to utilize pre-trained LLMs to solve EQA tasks without any additional fine-tuning Mu et al. (2024); Xiang et al. (2024); Huang et al. (2024). NaviLLM employed a scheme-based instruction that flexibly casts various tasks into generation problems, including the EQA task Zheng et al. (2024). OpenEQA employed a Frontier-Based Exploration (FBE) strategy for indoor environment exploration and tested the performance of various VLMs on the answer generation Majumdar et al. (2024). Besides, VLMs was also used to determine which room to explore in indoor environment based their commonsense reasoning capabilities Yin et al. (2025).
These agents, however, cannot be directly used for CityEQA tasks. Unlike indoor spaces, which are confined and divided into rooms, city spaces are vast and open. Agents in cities must navigate using landmarks and spatial relationships for long-term exploration Liu et al. (2024b). The proposed PMA addresses this by breaking down and planning for long-horizon CityEQA tasks, using large models across multiple modules to effectively handle open-ended questions and unseen environments.
@@ A.1 Dataset Collection and Validation
Any tasks identified as problematic were removed. To ensure the annotation consistency in the data collection and validation, we conducted Kappa statistical analyses for the raw annotation data from both Question and Answer revision phase and the task validation phase. The Kappa coefficients $\kappa$ for the two phases were 0.93 and 0.89, respectively, indicating a high level of agreement among annotators.
@@ A.3 Experiments Details
Figure 9: Categroy-level performance of the proposed PMA.
Figure 10: Examples of different EQA methods.
@@
  • Blind Agents. We choose four State-of-the-Art LLMs as blind agents, including GPT-4, Qwen-2.5, LLaMA-v3.1-8b, and DeepSeek-v3. They generate the answer purely based on the question, formulated as $\hat{y}=\text{LLMs}(q)$.
  • Socratic Agents. We sample efficient trajectories generated by H-EQA to simulate the observations available to Socratic Agents. Specifically, we select the last five frames from each trajectory and use GPT-4o to generate image captions $C$. Different LLMs—including GPT-4, Qwen-2.5, LLaMA-v3.1-8b, and DeepSeek-v3—are then used to produce the final answers, formulated as $\hat{y}=\text{LLMs}(q,C)$.
  • VQA Agents. They have direct access to images containing the answers. We use GPT-4o, Qwen-2.5, and LLaVA-v1.5-7b as VLMs to generate answers based on the images and questions, formulated as $\hat{y}=\text{VLMs}(q,p^{obs})$.
  • Exploring Agents. They are guided by different exploration strategies such as RE and FBE, and generate the answer based on the visual input $I$ at the termination position, formulated as $\hat{y}=\text{VLMs}(q,I)$. RE randomly selects an action from {MoveForward, TurnLeft, TurnRight, Stop} at each step. The angles for TurnLeft and TurnRight are set at 30°, and the distance for MoveForward is 10 meters, consistent with the setting of the Navigator in the PMA. FBE identifies the frontiers between explored and unexplored regions, samples one as the navigation point, and employs the A* algorithm to find a path. The maximum path length is also limited as 10 meters. To avoid excessive exploration, GPT-4o is employed to decide when to stop.
@@
Categroy-level performance of the PMA
The category-level performance of the proposed PMA is shown in Figure 9, and it varies across task types. PMA achieves the highest QAA on World Knowledge tasks, likely because these tasks rely partially on the LLM’s inherent knowledge and require minimal visual inputs. However, it performs the worst on Object Recognition tasks due to their open-ended answers and greater reliance on visual inputs.
Comparison between different EQA methods
We present the trajectories of PMA, H-EQA, and FBE to illustrate the different strategies adopted by them when searching for the answer to the same question, as shown in Figure 10. PMA finds the answer by decomposing the perception process into several sub-tasks and completing them step by step. H-EQA, with its stronger visual understanding and spatial reasoning abilities, can locate the answer in fewer steps. Moreover, H-EQA is often able to determine the answer from a greater distance, likely due to its extensive world knowledge, which allows it to fill in missing information even with incomplete observations. In contrast, FBE, lacking the ability to utilize landmarks such as building_1 and shop_1, can only fully explore the environment, resulting in lower perception efficiency. This highlights the differences between performing EQA tasks in urban spaces versus indoor environments.
Figure 11: Examples of the H-VQA and H-EQA.
Comparison among Human Agents
In Figure 11, we provide a case to illustrate why the performance of H-EQA is superior to that of HVQA. The given question is "What is the color of the car next to the red car?" The ground truth answer is "Black". HVQA was provided with the RGB image on the left for question answering. However, in this image, due to the influence of outdoor lighting, the originally black car appears gray, thus H-VQA provided an incorrect answer. In contrast, H-EQA can actively adjust the observation pose, observing the side of the car to reduce the impact of the lighting, and thereby providing the correct answer.
Analysis of Collector’s action
The statistics of various actions taken by Collector are shown in Figure 12. Besides, we present two cases to illustrate the effect of the collector. In the first case, as shown in Figure 13 (a), since the shop with black signboard was discovered too early in the Exploration stage, the starting pose of the collector was far from the target pose. Even after moving 10 steps promptly, it still failed to recognize the text on the black signboard. In the second case, as shown in Figure 13 (b), the yellow signboard that the collector needed to recognize was on the left side of the picture and seemed not to be fully displayed. At this time, the collector took the TurnLeft action, thus observing the entire yellow signboard and easily providing the correct answer.
@@
Figure 13: Examples of the Collection phase.